# Fast, accurate training and sampling of Restricted Boltzmann Machines

### v1

@@ Abstract
Thanks to their simple architecture, Restricted Boltzmann Machines (RBMs) are powerful tools for modeling complex systems and extracting interpretable insights from data. However, training RBMs, as other energy-based models, on highly structured data poses a major challenge, as effective training relies on mixing the Markov chain Monte Carlo simulations used to estimate the gradient. This process is often hindered by multiple second-order phase transitions and the associated critical slowdown. In this paper, we present an innovative method in which the principal directions of the dataset are integrated into a low-rank RBM through a convex optimization procedure. This approach enables efficient sampling of the equilibrium measure via a static Monte Carlo process. By starting the standard training process with a model that already accurately represents the main modes of the data, we bypass the initial phase transitions. Our results show that this strategy successfully trains RBMs to capture the full diversity of data in datasets where previous methods fail. Furthermore, we use the training trajectories to propose a new sampling method, parallel trajectory tempering, which allows us to sample the equilibrium measure of the trained model much faster than previous optimized MCMC approaches and obtain a better estimation of the log-likelihood. We illustrate the success of the training method on several highly structured datasets.
@@ 1 Introduction
Energy-based models (EBMs) are a classic approach to generative modeling that has been studied for decades. They were introduced using the Restricted Boltzmann Machine formulation by Smolensky [1] and later further developed by Sejnowski et al. [2]. They provide a straightforward method for modeling effective interactions within complex data distributions and for sufficiently simple energy functions, such as the Boltzmann machine (BM)  [3], it is also possible to interpret and infer the underlying constituent rules from the observed data. This inference strategy is often associated with the inverse Ising problem and pairwise interaction models [4], and it has found a great variety of applications in fields such as neuroscience [5] or computational biology [6]. A recent work has proposed replacing the use of pairwise models with the Restricted Boltzmann Machine (RBM) [7], as it allows the same direct interpretation of its energy function as an explicit many-body interaction model while greatly extending the expressive power of the model. RBMs are also very useful for grouping data into hierarchical families [8]. On the diametrically opposite side (on interpretability) are generative ConvNets [9, 10], where the energy function is formulated as a deep neural network, which are capable of synthesizing photorealistic images but are almost impossible to interpret as a physical model.
The applications of simple EBMs in science are very diverse. For example, they are often used today to encode the Hamiltonian of physical many-body systems, such as Quantum wave functions [11] or the accurate determination of ground state wave functions of strongly interacting and entangled quantum spins [12] or they have proven to be suitable for the representation of the AdS/CFT correspondence in theories of quantum gravity [13, 14]. Simple EBMs are also very common to encode the evolutionary constraints in protein families [6, 15], and to predict mutations [16], or to generate realistic synthetic sequences, such as fake human genomes [17, 18]. These examples show that, despite their somewhat old-fashioned architecture, shallow EBMs are increasingly seen as useful tools for better understanding modern physics/biology, as they allow for a certain level of analytical description.
@@
Despite the appealing modeling properties of RBMs, they are notoriously difficult to train, a challenge common to EBMs in general. The main difficulty arises from the computation of the log-likelihood gradient, which requires an ergodic exploration of a dynamically evolving and potentially complex free energy landscape using Markov Chain Monte Carlo (MCMC) processes. Recent studies have shown that models trained with non-convergent MCMC processes suffer from out-of-equilibrium dynamic memory effects [19, 20, 21]. This dynamical behavior can be explained analytically using moment-matching arguments  [19, 22]. While exploiting these effects can yield fast and accurate generative models, even for highly structured data [23] or high-quality images with RBMs [24], this approach results in a sharp separation between the model’s Gibbs-Boltzmann distribution and the dataset distribution, thereby undermining the interpretability of the model parameters  [22, 7]. Thus, to extract meaningful information from datasets using RBMs, it is essential to ensure proper mixing of the chains during training, in short, one needs equilibrium models.
@@
Both the ability to train an RBM in equilibrium and to generate convincing new samples from its equilibrium measure strongly depend on the dataset in question. For typical image datasets such as MNIST or CIFAR-10, good RBMs can be obtained by increasing the number of MCMC steps. However, this approach is no longer feasible for highly structured datasets [25]. Datasets from which one seeks scientific insights are often highly structured, such as genomics/proteomics data or low-temperature many-body physical systems. These datasets typically exhibit distinct clusters, identifiable via principal component analysis (PCA) which form distant groups of similar entries. We show an example of the PCA of 3 clustered dataset we will be studying in this work in Fig. 1; details about these datasets are given in the caption and in the Supplemental Information (SI). During training, the model must evolve from an initial normal distribution to an increasingly multimodal distribution. Sampling from multimodal distributions is particularly challenging because the mixing times are determined by the transition times between modes. But this is not the only difficulty. These distant modes are encoded by second-order phase transitions during training [26, 27, 28], leading to diverging mixing times in these regions — a phenomenon known as critical slowdown —, which means that mixing times are expected to grow with a power of their system size. This sampling challenge not only hinders the training process, but also limits the model’s ability to generate new samples. Obtaining new and independent configurations would require an impractically large number of sampling steps.
Figure 1: Clustered datasets. In A-C we show the 4 different clustered data sets that we will consider in this paper, projected onto their first two PCA components. In A we show the data of the MNIST 01 dataset (both projected and some instances), which contains only the 0-1 images of the complete MNIST dataset. In B, we show the Mickey dataset, an artificial dataset whose PCA forms a “Mickey"” face shape. In C, we show data from the Human Genome Dataset (HGD), which contains binary vectors each corresponding to a human individual and whose sites correspond to selected genes. A value of 1 at a particular position means that a mutation was observed there compared to an individual reference sequence. Details of these data sets can be found in the SI. In D-F we show the samples we generate with the low-rank RBMs that are used as initial point of a standard training.
@@ 2 Related work
Training EBMs by maximizing log-likelihood has long been a challenge in the community [29, 10]. EBMs gained popularity with the introduction of the contrastive divergence algorithm [30], in which a set of parallel chains is initialized on independent examples in the minibatch and the MCMC process iterates for a few steps. Despite its widespread use, this algorithm yields models with poor equilibrium properties that are ineffective as generative models [31, 32, 21]. An improvement is the persistent contrastive divergence (PCD) algorithm [33], which maintains a permanent chain in which the last configurations used to estimate the previous gradient update are reused. PCD acts like a slow annealing process improving gradient estimation quality. However, it often fails on clustered data as the statistical properties of the permanent chain quickly move away from the equilibrium measure and degrade the model [25]. This problem, which is primarily related to phase coexistence, can be addressed with constrained MCMC methods if appropriate order parameters are identified. For RBMs, these order parameters are related to the singular value decomposition of the model coupling matrix, which enables efficient reconstruction of multimodal distributions [25]. Although this method is effective for evaluating model quality, it is too computationally intensive to be used in training, even if it leads to models with good equilibrium properties. Other optimized MCMC methods, such as the Parallel Tempering (PT) [34] algorithm, simulate multiple models at different temperatures, facilitating mixing through temperature exchange [35, 32]. However, PT is costly and often ineffective, especially because EBMs undergo first-order phase transitions at the temperature where PT typically fails because one needs too many temperatures to make the moves accepted. We will see below that a more appropriate approach exchanges the models at different training times, which only implies crossing second-order phase transitions.
The population annealing algorithm, which reweights parallel chains during learning based on their relative weight changes during parameter updates, was proposed as an alternative [36]. Similarly, reweighting chains using non-equilibrium physics concepts such as the Jarzynski equality has been proposed [37]. Both approaches struggle with highly structured data sets. To prevent the different chains to get too correlated around the training phase transitions, one must either increase the number of sampling steps or decrease the learning rate, which in practice means very long training processes to ensure a proper equilibrium training. Another strategy is to use EBMs as corrections for straightforward-to-sample flow-based models [38]. This simplifies sampling and learning, but sacrifices the interpretability of the energy function, which was our goal. An evolving flow model can be used as a fast sampling moves proposer for the EBM [39] objective. This method requires the training of two different networks in parallel and may result in the drop of the move acceptancy as the EBM becomes specialized.
@@
For RBMs, a recent method called “stacked tempering" [40] dramatically speeds up sampling by training smaller RBMs with latent variables from previous models, allowing fast updates to be proposed using a PT like algorithm. Authors also showed that this algorithm was much faster than the standard PT. While effective, it is too cumbersome for use in training. Also for RBMs, it has recently been shown that it is possible to train a low-rank RBM that accurately reproduces the statistics of the data projected along the $d$ first data principal directions through a convex and very fast optimization process (see [41] and the discussion below). This low-rank model can be seen as a good approximation to the correct RBM needed to describe the data, and has the nice property that it can be efficiently sampled via a static Monte Carlo process.
In this paper, we will show how to drastically reduce training times by starting the RBM training process at this low-rank RBM, as this means that the first and strongest dynamic effects associated with them are directly bypassed. We also show that one can exploit the training trajectory to develop an effective sampling method, the parallel trajectory tempering (PTT) that outperforms the “stacked tempering" [40] and only requires saving a reduced number of models during the training. This strategy also allows to obtain reliable estimations for the log-likelihood in well-trained models, much better than those obtained with the standard Annealing Important Sampling (AIS) techniques [42]. Using both strategies, we show that we are able to train and evaluate methods that accurately represent the different modes in the dataset, where standard methods lead to mode collapse effects.
@@ 3 The Restricted Boltzmann Machine
where $Z$ is the partition function of the system. As with other models containing hidden variables, the training objective is to minimize the distance between the empirical distribution of the data, $p_{\mathcal{D}}(\bm{v})$, and the model’s marginal distribution over the visible variables, $p(\bm{v})=\sum_{\bm{h}}\exp\left[-\mathcal{H}(\bm{v},\bm{h})\right]/Z=\exp\left[-H(\bm{v})\right]/Z$. Minimizing the Kullback-Leibler divergence is equivalent to maximizing the likelihood of observing the dataset in the model. Thus, the log-likelihood $\mathcal{L}=\langle-H(\bm{v})\rangle_{\mathcal{D}}-\log Z$ can be maximized using the classical stochastic gradient ascent. For a training dataset $\mathcal{D}=\{{\bm{v}}^{(m)}\}_{m=1,\dots,M}$, the log-likelihood gradient is given by
@@
where $\left\langle{\cdot}\right\rangle_{\mathcal{D}}$ denotes the average with respect to the entries in the dataset, and $\left\langle{\cdot}\right\rangle_{\mathrm{RBM}}$ with respect to $p(\bm{v},\bm{h})$. Since $Z$ is intractable, the model averages in the gradient are typically estimated using $N_{\mathrm{s}}$ independent MCMC processes, and observable averages $\left\langle{o(\bm{v},\bm{h})}\right\rangle_{\mathrm{RBM}}$ are replaced by $\sum_{r=1}^{R}o(\bm{v}^{(r)},\bm{h}^{(r)})/R$, with $(\bm{v}^{(r)},\bm{h}^{(r)})$ being the last configurations reached with each of the $r=1,\ldots,R$ parallel chains. To obtain reliable estimates, it should be ensured that each of the Markov chains mix well before each parameter update. However, ensuring equilibrium at each update is impractical, slow and tedious. The common use of non-convergent MCMC processes is the cause of most difficulties and weird dynamical behaviors encountered in training RBMs [21].
Typical MCMC mixing times in RBMs are very small at the beginning of the training and grow as it progresses [21], suffering with sharp increases every-time the training trajectory crosses each of the critical transitions that give birth to new modes [28]. In order to minimize out-of-equilibrium effects, it is often useful to keep $R$ permanent (or persistent) chains, which means that the last configurations reached with the MCMC process used to estimate the gradient at a given parameter update $t$, $\bm{P}_{t}\equiv\{(\bm{v}^{(r)}_{t},\bm{h}^{(r)}_{t})\}_{r=1}^{R}$, are used to initialize the chains of the subsequent update $t+1$. This algorithm is typically referred as PCD. In this scheme, the process of training can be mimicked to a slow cooling process, only that instead of varying a single parameter, the temperature, a whole set of parameters $\bm{\Theta}_{t}=(\bm{w}_{t},\bm{\theta}_{t},\bm{\eta}_{t})$ are updated at every step to $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}+\gamma\bm{\nabla}\mathcal{L}_{t}$ with $\bm{\nabla}\mathcal{L}_{t}$ being the gradient in Eq. (2) estimated using the configurations in $\bm{P}_{t}$, and $\gamma$ being the learning rate.

### v2

# Fast training and sampling of Restricted Boltzmann Machines

@@ Abstract
Restricted Boltzmann Machines (RBMs) are effective tools for modeling complex systems and deriving insights from data. However, training these models with highly structured data presents significant challenges due to the slow mixing characteristics of Markov Chain Monte Carlo (MCMC) processes. In this study, we build upon recent theoretical advancements in RBM training, focusing on the gradual encoding of data patterns into singular vectors of the coupling matrix, to significantly reduce the computational cost of training (in very clustered datasets) and evaluating and sampling in RBMs in general. The learning process is analogous to thermodynamic continuous phase transitions observed in ferromagnetic models, where new modes in the probability measure emerge in a continuous manner. Such continuous transitions are associated with the critical slowdown effect, which adversely affects the accuracy of gradient estimates, particularly during the initial stages of training with clustered data. To mitigate this issue, we propose a pre-training phase that encodes the principal components into a low-rank RBM through a convex optimization process. This approach facilitates efficient static Monte Carlo sampling and accurate computation of the partition function. Furthermore, we exploit the continuous and smooth nature of the parameter annealing trajectory to achieve reliable and computationally efficient log-likelihood estimations, enabling online assessment during the training process, and proposing a novel sampling strategy termed parallel trajectory tempering that outperforms previously optimized MCMC methods. Our results demonstrate that this innovative training strategy enables RBMs to effectively address highly structured datasets that conventional methods struggle with. Additionally, we provide evidence that our log-likelihood estimation is more accurate than traditional, more computationally intensive approaches in controlled scenarios. Moreover, the parallel trajectory tempering algorithm significantly accelerates MCMC processes compared to existing and conventional methods.
@@ 1 Introduction
Energy-based models (EBMs) are a long-established approach to generative modeling, with the Boltzmann Machine (BM)  (Hinton & Sejnowski, 1983) and Restricted Boltzmann Machine (RBM) (Smolensky, 1986; Ackley et al., 1985) being among the earliest examples. They offer a clear framework for capturing complex data interactions and, with simple energy functions, can reveal underlying patterns, making them especially valuable for scientific applications. Pairwise-interacting models (Nguyen et al., 2017) have been widely used for decades for inference applications in fields like neuroscience (Hertz et al., 2011) and computational biology (Cocco et al., 2018). More recent research has shown that RBMs can also be reinterpreted as many-body physical models (Decelle et al., 2024; Bulso & Roudi, 2021), analyzed to identify hierarchical relationships in the data (Decelle et al., 2023), or used to uncover biologically relevant constituent features (Tubiana et al., 2019; Fernandez-de Cossio-Diaz, 2025). In essence, RBMs present a great compromise between expressibility and interpretability: they are universal approximators capable of modeling complex datasets and simple enough to allow a direct extraction of insights from data. This interpretative task is particularly challenging when dealing with more powerful models such as generative convolutional networks EBMs (Xie et al., 2016; Song & Kingma, 2021) and other state-of-the-art models in general. At this point, it is also important to stress that RBM are not only interpretable, but they perform particularly well when dealing with tabular datasets with discrete variables, and where data is not abundant, which is often the case in genomics/proteomics or neural recording datasets. For instance, RBMs have shown to be among the most reliable approaches to generate synthetic human genomes (Yelmen et al., 2021; 2023) or to model inmunological response (Bravi et al., 2021a; b; 2023) or neural activity (van der Plas et al., 2023; Quiroz Monnens et al., 2024).
@@
Despite the appealing modeling properties of RBMs, they are notoriously difficult to train, a challenge common to EBMs in general. The main difficulty arises from the computation of the log-likelihood gradient, which requires an ergodic exploration of a dynamically evolving and potentially complex free energy landscape using Markov Chain Monte Carlo (MCMC) processes. Recent studies have shown that models trained with non-convergent MCMC processes suffer from out-of-equilibrium dynamic memory effects (Nijkamp et al., 2019; 2020; Decelle et al., 2021). This dynamical behavior can be explained analytically using moment-matching arguments  (Nijkamp et al., 2019; Agoritsas et al., 2023). While exploiting these effects can yield fast and accurate generative models, even for highly structured data (Carbone et al., 2023) or high-quality images with RBMs (Liao et al., 2022), this approach results in a sharp separation between the model’s Gibbs-Boltzmann distribution and the dataset distribution, thereby undermining the interpretability of the model parameters as investigated in Agoritsas et al. (2023); Decelle et al. (2024). Thus, to extract meaningful information from datasets using RBMs, it is essential to ensure proper mixing of the chains during training, in short, one needs equilibrium models.
@@
Both the ability to train an RBM in equilibrium and to generate convincing new samples from its equilibrium measure strongly depend on the dataset in question. For typical image datasets such as MNIST or CIFAR-10, good RBMs can be obtained by increasing the number of MCMC steps. However, this approach is no longer feasible for highly structured datasets (Béreux et al., 2023). Natural datasets from which one seeks scientific insights are often highly structured, such as genomics/proteomics data, neural-recordings or low-temperature configurations in physical systems. These datasets typically exhibit distinct clusters, identifiable via principal component analysis (PCA) which form distant groups of similar entries often associated to different labels. We show an example of the PCA of the 5 datasets we will be studying in this work in Fig. 1. The first 4 datasets are very clustered, while CelebA is not, as it presents a rather compact PCA. A very similar (non-clustered) situation is obtained if the entire MNIST dataset is considered. Details about these datasets are given in the caption of figure 1. During training, the model must evolve from an initial normal distribution to an increasingly multimodal distribution. Sampling from multimodal distributions is particularly challenging because the mixing times are controlled by the transition times between modes. But this is not the only difficulty. In RBMs, these distant modes can be shown analytically(Decelle et al., 2017; 2018; Bachtis et al., 2024) to gradually appear in the model’s equilibrium measure during training by a process formally analogous to the ferromagnetic phase transition in simple spin models, the process by which the magnetic spins of a material spontaneously align with a fixed pattern when the temperature goes below a certain transition temperature. The RBM training is characterized as a sequence of such phase transitions in which the specialization of the model parameters plays the role of temperature reduction. These pattern-encoding phase transitions are continuous phase transitions (also known as second-order transitions or critical transitions), and these types of transitions are associated with a sudden arrest in the characteristic times of the visible variables, a phenomenon known as critical slowdown, which causes the MCMC mixing time to diverge with a power of the system size when the transition point is approached. The critical slowdown effects that occur in the training trajectory were recently studied in Bachtis et al. (2024). A brief introduction to phase transitions and the associated dynamical effects can be found in the SI G.
This sampling challenge not only hinders the training process, but also limits the model’s ability to generate new samples. Obtaining new and independent configurations would require an impractically large number of sampling steps.
Figure 1: Datasets. Panels A-E display 5 distinct datasets projected onto their first two PCA components. In some instances, the dots are color-coded to indicate different labels. In A, the MNIST 01 dataset, featuring images of the digits 0 and 1 from the complete MNIST collection, along with a few sample images. In B, the “Mickey” dataset, an artificial dataset whose PCA representation forms the shape of Mickey Mouse’s face. In C, the Human Genome Dataset (HGD), which consists of binary vectors representing mutations or non-mutations for individuals compared to a reference genome across selected genes. In D, the Ising dataset, showcasing equilibrium configurations of the 2D ferromagnetic Ising model at an inverse temperature of $\beta=0.44$. In E, the CelebA dataset in black and white, resized to 32x32 pixels. For more details on these datasets, please refer to the SI.
@@ 2 Related work
Training EBMs by maximizing log-likelihood has long been a challenge in the community (LeCun et al., 2006; Song & Kingma, 2021). EBMs gained popularity with the introduction of the contrastive divergence algorithm (Hinton, 2002), in which a set of parallel chains is initialized on independent examples in the minibatch and the MCMC process iterates for a few steps. Despite its widespread use, this algorithm yields models with poor equilibrium properties that are ineffective as generative models (Salakhutdinov & Murray, 2008; Desjardins et al., 2010; Decelle et al., 2021). An improvement is the persistent contrastive divergence (PCD) algorithm (Tieleman, 2008), which maintains a permanent chain in which the last configurations used to estimate the previous gradient update are reused. PCD acts like a slow annealing process improving gradient estimation quality. However, it often fails on clustered data as the statistical properties of the permanent chain quickly move away from the equilibrium measure and degrade the model (Béreux et al., 2023). This problem, which is primarily related to phase coexistence, can be addressed with constrained MCMC methods if appropriate order parameters are identified. For RBMs, these order parameters are related to the singular value decomposition of the model coupling matrix, which enables efficient reconstruction of multimodal distributions (Béreux et al., 2023). Although this method is effective for evaluating model quality, it is too computationally intensive to be used in training, even if it leads to models with good equilibrium properties. Other optimized MCMC methods, such as the Parallel Tempering (PT) algorithm  (Hukushima & Nemoto, 1996) (in which the model is simulated in parallel at different temperatures which are exchanged once in a while using the Metropolis rule), have been proposed to improve the training process (Salakhutdinov, 2009; Desjardins et al., 2010) and show important improvements in certain datasets (Krause et al., 2020).
However, PT is expensive because it requires sampling at multiple temperatures while using samples from only one model. It’s also often ineffective with highly clustered data due to first-order phase transitions in EBMs, where modes disappear abruptly at certain temperatures, as discussed by Decelle & Furtlehner (2021a). This leads to sharp drops in acceptance rates for exchanges between temperatures near the transition. For this reason, first-order transitions typically hinder PT, as many intermediate temperatures are needed to allow sufficient acceptance rates in the exchanges. In contrast, continuous transitions, as seen during the training trajectory, cause modes to gradually separate as the model parameters are changed.
The population annealing algorithm, which adjusts the statistical weights of parallel chains during learning each time the parameters are updated, has been proposed as an alternative approach (Krause et al., 2018). Reweighting chains using non-equilibrium physics concepts like the Jarzynski equation has also been explored (Carbone et al., 2024). However, both methods struggle with highly structured data, as frequent reweighting (which involves duplicating only a few samples) biases the equilibrium measure due to an insufficient number of effective independent chains. To prevent chains from becoming too correlated around transition regions during training, either more sampling steps or a lower learning rate is required, which leads to very long training times to ensure equilibrium. In our case, we tried to used both methods with our datasets but the training becomes too unstable or two slow if learning rate and number of Gibbs steps are auto adjusted. Another approach uses EBMs as corrections for flow-based models that are easier to sample (Nijkamp et al., 2022), simplifying learning but sacrificing the interpretability of the energy function, a key goal of this work. Alternatively, evolving flow models can serve as fast sampling proposers for the EBM (Grenioux et al., 2023), but this requires training two separate networks and the acceptance of the proposed moves is significantly reduced as the EBM becomes more specialized.
@@
A recent method for RBMs, called “stacked tempering” (Roussel et al., 2023), significantly accelerates sampling by training smaller RBMs using latent variables from previous models. This allows for fast updates through a PT-like algorithm, and the authors demonstrated that it is much faster than standard PT. However, while this method is effective for sampling a trained model, it is not suitable for learning, as it requires sequentially training multiple models stacked on top of each other. Also for RBMs, it has recently been shown in a theoretical work that it should be possible to train a low-rank RBM that accurately reproduces the statistics of the data projected along the $d$ first data principal directions through a convex and very fast optimization process (see Decelle & Furtlehner (2021a) and the discussion below) using a mapping to another model, the so-called Restricted Coulomb Machine. This low-rank model can be seen as a good approximation to the correct RBM needed to describe the data, and has the nice property that it can be efficiently sampled via a static Monte Carlo process. While this work showed the efficiency of this training process to describe simple synthetic datasets embedded in only one or two dimensions, the use of this algorithm to obtain reliable low-rank RBMs with real data requires further developments that will be discussed in this paper, in particular: (i) the training of a bias, which is crucial to gain almost for free an additional dimension, and to deal with image datasets, and (ii) a tuning of the RBM fast sampling process, which was not properly investigated in the original paper.
In this paper, we show how starting the RBM training with a low-rank RBM in very clustered datasets can drastically improve the quality of the models by circumventing the initial dynamical arrest effects that occur during the training process, at a very similar computational cost. We also show that the training trajectory can be leveraged to estimate the log-likelihood at a very little computational effort, so that both the training and test set log-likelihood can be easily and reliably computed online during training. Furthermore, we propose an efficient sampling method, which we call Parallel Trajectory Tempering (PTT). This method outperforms optimized MCMC methods such as the traditional Parallel Tempering (PT) or the “Stacked Tempering” (Roussel et al., 2023) only requiring storing the parameters of the model at a few stages during training. By combining these strategies, we are able to train equilibrium models that accurately capture different modes in datasets prone to mode collapse using standard methods, easily evaluate RBMs during the training process, and easily sample from multimodal distribution.
@@ 3 The Restricted Boltzmann Machine
where $Z$ is the partition function of the system. As with other models containing hidden variables, the training objective is to minimize the distance between the empirical distribution of the data, $p_{\mathcal{D}}(\bm{v})$, and the model’s marginal distribution over the visible variables, $p(\bm{v})=\sum_{\bm{h}}\exp\left[-\mathcal{H}(\bm{v},\bm{h})\right]/Z=\exp\left[-H(\bm{v})\right]/Z$. Minimizing the Kullback-Leibler divergence is equivalent to maximizing the likelihood of observing the training dataset in the model. Thus, the log-likelihood $\mathcal{L}=\langle-H(\bm{v})\rangle_{\mathcal{D}}-\log Z$ can be maximized using the classical stochastic gradient ascent. For a training dataset $\mathcal{D}=\{{\bm{v}}^{(m)}\}_{m=1,\dots,M}$, the log-likelihood gradient is given by
@@
where $\left\langle{\cdot}\right\rangle_{\mathcal{D}}$ denotes the average with respect to the entries in the dataset, and $\left\langle{\cdot}\right\rangle_{\mathrm{RBM}}$ with respect to $p(\bm{v},\bm{h})$. Since $Z$ is intractable, the model averages in the gradient are typically estimated using $N_{\mathrm{s}}$ independent MCMC processes, and observable averages $\left\langle{o(\bm{v},\bm{h})}\right\rangle_{\mathrm{RBM}}$ are replaced by $\sum_{r=1}^{R}o(\bm{v}^{(r)},\bm{h}^{(r)})/R$, with $(\bm{v}^{(r)},\bm{h}^{(r)})$ being the last configurations reached with each of the $r=1,\ldots,R$ parallel chains. To obtain reliable estimates, it should be ensured that each of the Markov chains mix well before each parameter update. However, ensuring equilibrium at each update is impractical, slow and tedious. The common use of non-convergent MCMC processes is the cause of most difficulties and weird dynamical behaviors encountered in training RBMs as discussed in Decelle et al. (2021). In order to minimize out-of-equilibrium effects, it is often useful to keep $R$ permanent (or persistent) chains, which means that the last configurations reached with the MCMC process used to estimate the gradient at a given parameter update $t$, $\bm{P}_{t}\equiv\{(\bm{v}^{(r)}_{t},\bm{h}^{(r)}_{t})\}_{r=1}^{R}$, are used to initialize the chains of the subsequent update $t+1$. This algorithm is typically referred as PCD. In this scheme, the process of training can be mimicked to a slow cooling process, only that instead of varying a single parameter, e.g. the temperature, a whole set of parameters $\bm{\Theta}_{t}=(\bm{w}_{t},\bm{\theta}_{t},\bm{\eta}_{t})$ are updated at every step to $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}+\gamma\bm{\nabla}\mathcal{L}_{t}$ with $\bm{\nabla}\mathcal{L}_{t}$ being the gradient in equation 2 estimated using the configurations in $\bm{P}_{t}$, and $\gamma$ being the learning rate.
Typical MCMC mixing times in RBMs are short at the beginning of training but increase as it progresses (Decelle et al., 2021). Recent analytical and numerical studies show that the mixing times exhibit sharp, localized peaks when the model encodes a new mode during training (Bachtis et al., 2024). This phenomenon, known in physics as critical slowing down, is related to the system going through critical phase transitions. For RBMs, the first phase transitions during maximum likelihood training are associated with the encoding of the principal components of the dataset into the singular vectors of the weight matrix $\bm{W}$ (Decelle et al., 2017; 2018; Decelle & Furtlehner, 2021b; Bachtis et al., 2024). Since these directions can be determined before training by performing PCA on the data, one can pre-train an RBM to encode them accurately. This avoids crossing the initial phase transitions encountered during the standard training and the associated degradation of the gradient calculation.